Based on the time series values from problem number 2, consider the following table of exponential smoothing values using ? = ?. ? for the time series.

Month

Units Sold (Thousands)

Forecast (F)

error

Squared error

1

9

*

* *

2

3

(i)?

(ii)?

36

3

6

7.2000

-1.2000

1.44

4

6

6.8400

-0.8400

0.7056

5

12

6.5880

5.4120

29.2897

6

9

8.2116

0.7884 0.6216

7

(iii)?

a) b) c) d) e)

a. Compute the number (i): Show your work for full credit
b. Compute the number (ii): Show your work for full credit
c. Compute the number (iii): Show your work for full credit
d. What is the mean square error (MSE)? Show your work for full credit
Answer #1

a) Assuming Forecast for first month = Units sold = 9

Forecast for Month 2: F2 = F1 + ?*(A1-F1) = 9+0.3*(9-9) = 9

b) Error for month 2 = Actual-Forecast = 3-9 = -6

c) Forecast for Month 7: F7 = F6 + ?*(A6-F6) = 8.2116+0.3*(9-8.2116) = 8.2116+0.3*0.7884 = 8.2116+0.2365 = 8.4481

d) MSE = Sum of squared errors/n = (36+1.44+0.7056+29.2897+0.6216)/5 = 68.0569/5 = 13.6114